OpenStudy (anonymous):

find a parabola y=ax^2 + bx + c that passes through the point ( 1, 4 ) and whose tangent lines at x=-1 and x=5 have slopes of 6 and -2 respectively.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

we gotta make some equations

OpenStudy (misty1212):

first one says \[4=a+b+c\]

OpenStudy (misty1212):

the derivative is \[2ax+b\] so the second one says \[-2a+b=6\]

OpenStudy (misty1212):

third one says \[10a+b=-2\]

OpenStudy (misty1212):

solve the system \[-2a+b=6\\ 10a+b=-2\] for \(a\) and\(b\), then replace in the first one to find \(c\)

OpenStudy (anonymous):

how did you get 10a + b = -2?